You have been asked to mix up some 70% isopropyl alcohol solution. You have 95% isopropyl alcohol solution on hand. How much do you have to add to 750 mL of pure water to created a 70% solution?
Answer by stanbon(75887)

You have been asked to mix up some 70% isopropyl alcohol solution. You have 95% isopropyl alcohol solution on hand. How much do you have to add to 750 mL of pure water to created a 70% solution?
-------
Equation:
iso + iso = iso
0.95x + 0*750 = 0.70(x+750)
---
Multiply thru by 100 to get:
95x + 0 = 70x + 70*750
25x = 70*750
x = 70*30
x = 2100 mL (amt. of 95% iso to be added)
=============================================
Cheers,
Stan H.
==================